OpenStudy (anonymous):

2 sin 3theta -sqrt2 = 0

OpenStudy (anonymous):

\[2\sin 3\theta-\sqrt{2}=0=>2\sin 3\theta=\sqrt{2}=>\sin 3\theta=\frac{ 1 }{ \sqrt{2} }=>3\theta=\sin^{-1} \left( \frac{ 1 }{ \sqrt{2} } \right)=>3\theta=45=>\theta=15\]

OpenStudy (anonymous):

2 sin 3theta -sqrt2 = 0 2 sin(3theta)=sqrt(2) sin(3 theta) = sqrt(2)/2 3theta = arcsin(sqrt(2)/2) theta = arcsin(sqrt(2)/2)/3 theta = (PI/4 + 2PI*n)/3 where n belongs to Z theta = PI/12 + 2PI*n/3 where n belongs to Z

OpenStudy (anonymous):

\[3\theta=45=>\theta=15 (answer)\]

OpenStudy (anonymous):

you typed it in degrees however sin has a period of 2PI thus there is an infinite set of answers

OpenStudy (anonymous):

but here its not mentioned that in which form find the answer..!!!if there is any restricton then it colud sooo..1!!

OpenStudy (anonymous):

Exactly. No bounds specified then you need to show all possible answers.

OpenStudy (anonymous):

okay..!!